# Chi-Hsi Kung

### v1

@@ 1 Introduction
We evaluate the learned procedure-aware video representations in three key procedural video understanding tasks—error detection, temporal action segmentation, and action phase classification and frame retrieval—and show that they significantly enhance performance compared to strong baselines. Our contributions are summarized as follows:
@@ 3 Pretraining Objective:
Video-Level Alignment The goal of this loss is to align video-level visual embeddings to summary text embeddings and to enhance procedural awareness using video-level counterfactuals. We first obtain video-level visual embeddings from clip-level embeddings by leveraging a self-attention block as an aggregator function. Then, using contrastive learning, each visual embedding is aligned to its corresponding summary text embedding and contrasted against text embeddings of the generated counterfactual Misordered and Missing-step. The formulation is defined as,
@@ 4 Experiments
Tasks We evaluate our model’s learned representations on four procedural video-understanding tasks: temporal action segmentation on GTEA [5] and EgoPER [12], error detection on EgoPER [12], and action phase classification and frame retrieval on the Align-Ego-Exo dataset [28]. We report F1 score, mean average precision (mAP), edit distance, frame accuracy, or Error Detection Accuracy (EDA) metrics based on each task convention.

### v2

@@ 1 Introduction
We evaluate the learned procedure-aware video representations in three key procedural video understanding tasks—error detection, temporal action segmentation, and action phase classification and frame retrieval—and show that they significantly enhance performance compared to strong baselines. In summary, our contributions are:
@@ 3 Pretraining Objective:
Video-Level Alignment This loss aligns video-level visual embeddings to summary text embeddings and enhances procedural awareness using video-level counterfactuals. We first obtain video-level visual embeddings from clip-level embeddings by using a self-attention block as an aggregator function. Then, using contrastive learning, each visual embedding is aligned to its corresponding summary text embedding and contrasted against text embeddings of the generated counterfactual Misordered and Missing-step,
@@ 4 Experiments
Tasks We evaluate our model’s learned representations on four procedural video-understanding tasks: temporal action segmentation on GTEA [5] and EgoPER [12], error detection on EgoPER [12], and action phase classification and frame retrieval on Align-Ego-Exo [28]. We report F1 score, mean average precision (mAP), edit distance, frame accuracy, or Error Detection Accuracy (EDA) metrics based on each task convention.
@@ 5 Conclusions
Acknowledgments. CK, FR, JH, DC were supported in part by NSF (DRL-2112635 via AI Institute for Engaged Learning) and Lilly Endowment via PTI. YC was supported in part by National Science and Technology Council (113-2628-E-A49-022, 114-2628-E-A49-007), NYCU Higher Education Sprout Project, and Ministry of Education Yushan Fellow Program Administrative Support Grant.